OpenStudy (anonymous):

1/p+1/q=1/f solve for p

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

q/pq +p/pq=1/f (P+q)/pq=1/f

OpenStudy (anonymous):

cross multiply

OpenStudy (anonymous):

p+q/pq=1/f p+q=pq/f p=(pq/f) - q

OpenStudy (anonymous):

1/p+1/f=1/q subtract 1/f from both sides 1/p = 1/q-1/f cross multiply right hand side 1/p = (q-f)/qf invert both sides (1/???) p=qf/(q-f)